# The Fast and the Frame-Dragging: Efficient waveforms for asymmetric-mass eccentric equatorial inspirals into rapidly-spinning black holes

### v1

@@ III.1.2 Rapid computation of ODE derivatives
In order to integrate Eqs. 2 and 3, we must first specify the form of $\hat{f}^{(0)}_{p,e}$ and $\hat{\Omega}_{\phi,r}$. The latter terms are simple to incorporate; analytic expressions for the dimensionless orbital frequencies of geodesics in generic Kerr, which are equal to $\hat{\Omega}_{\phi,r}$ at adiabatic order, have existed in the literature for some time (see e.g. Fujita:2009bp ; vandeMeent:2019cam and references therein) and are inexpensive to evaluate. We employ these expressions in our framework, simplifying terms involving $x_{I}$ (given $x_{I}=1$) where possible for efficiency and numerical stability.
Obtaining the forcing functions is less straightforward, as the data generation procedure in Section III.1.1 is orders of magnitude too expensive to be applied on-the-fly during trajectory integration. However, these fluxes vary smoothly over the parameter space and are ideal targets for interpolation schemes. We therefore opt to generate forcing function values en-masse, producing data grids in parameter space over which to perform this interpolation. Producing these data grids is computationally expensive, but must only be performed once; as the methods described in Section III.1.1 are straightforward to parallelise, this can be achieved relatively quickly on distributed computing resources. Of particular note is the accuracy to which the fluxes must be estimated, which is $\mathcal{O}(\nu)$ for sub-radian accumulated phase error over the course of a typical trajectory evaluation.
@@
Next, we must specify our interpolation variables. This is essential for our framework because a regular grid cannot be constructed in $(a,p,e)$ space (recalling that $x_{I}=1$ for equatorial inspirals) due to the boundary imposed by $p_{\mathrm{sep}}$ that varies across the parameter space. Choosing sensible interpolation variables can also spread out rapid variation in the fluxes across the parameter space (especially near the separatrix), improving the accuracy of interpolation in these regions. To this end, we format our data grids in terms of three interpolation variables $(u,w,z)$ that are related to $(a,p,e)$ by bijective and analytically-invertible transformations. For brevity, we will not state these transformations in the main text here (they are given in full in Appendix B) but instead summarize the important features of the resulting grid structure below:
@@
The shape and grid-point distribution of these grids is also shown in Fig. 20 and Fig. 21 in Appendix B, which clearly displays the structure of the eccentricity taper. Functions for performing these mappings are implemented in few and are readily applicable in the construction of user-specified grids.
Finally, we must choose the dimensions of the grid. With all other aspects of the problem fixed, this is essentially a trade-off between the accuracy of the interpolant and the computational cost of producing the grids. However, assessing how interpolation error in trajectory modelling propagates to the accuracy of the resulting waveform is not straightforward. We therefore opt to identify empirically what flux grid dimensions are required to satisfy waveform accuracy requirements. Based on this process, we settle on the dimensions $(N_{u},N_{w},N_{z})$ = $(129,65,65),(65,33,33)$ for the inner and outer grid regions respectively. As we will later demonstrate in Section IV, this ensures that grid point density (and therefore, the accuracy of the interpolated forcing functions) is sufficient for our model to be highly robust over the majority of the parameter space.
@@ III.1.3 ODE integration with continuous solution
A continuous approximation for the phase also yields similar piecewise polynomials for the orbital frequencies and their derivatives with respect to time; as shown in Fig. 2, this approximation is very accurate (exceeding 9 decimal places in the frequencies for the majority of the inspiral). This is particularly useful in the construction of frequency-domain waveforms (which require this information as an input) and will enable resonance effects (which require careful monitoring of these frequencies, as well as accurate frequency derivatives) to be seamlessly folded into the few framework in the future.
@@ III.1.4 Numerical integration of ODE system
With our framework for obtaining the derivatives in Eqs. 2 and 3 and integrating the resulting ODE system in place, we are ready to evolve a trajectory from a set of initial conditions $\{p_{0},e_{0},\Phi_{\phi 0},\Phi_{r0}\}$ given specific values for $M$, $\nu$ and $a$. To improve the numerical stability of the integrator, we first rescale the problem such that the integration variables vary by $\mathcal{O}(1)$ across an inspiral. We achieve this by integrating the system with respect to the radiation-reaction timescale $t_{\mathrm{rr}}=\nu t$ and eliminating the pre-factor $\nu$ from Eq. 2 accordingly. Rather than scaling Eq. 3 similarly, we leave it unchanged (in order to achieve the desired property of $\mathcal{O}(1)$ scaling for all parameters) and instead rescale the orbital phase evolution post-integration by $\nu^{-1}$. This approach is valid in the adiabatic approximation because (by definition) the $\hat{f}^{(0)}_{p,e}$ are independent of the orbital phases. A consequence of this choice is that, for an integrator absolute error tolerance $\sigma_{\mathrm{tol}}$, the effective error tolerance on the orbital phases is $\sigma_{\mathrm{tol}}/\nu$. This is in fact a desirable outcome: we are mainly interested in tuning the accuracy with which $p(t)$ and $e(t)$ are evolved (as small errors in these parameters will gradually accumulate over an inspiral, coupling in turn to the orbital phase evolution), whereas the error in the reconstructed phase will remain of order $\sigma_{\mathrm{tol}}/\nu$, which is small and has negligible impact on waveform phase accuracy. This scaling is demonstrated in Fig. 3, where phase error grows in proportion to (but is several orders of magnitude larger than) the error tolerance of the ODE integrator. Trajectory wall-time varies significantly with respect to ODE error; to ensure trajectories are accurate to within $10^{-3}$ radians, we set $\sigma_{\mathrm{tol}}=10^{-11}$ by default, but this can be relaxed by the end user according to their accuracy requirements. We also note that trajectory wall-time will be hardware-dependent; the timings presented in Fig. 3 were obtained with an Apple M3 processor.
@@
The size of the first integration step from the starting point is chosen automatically with standard techniques GLADWELL1987175 . The integrator then steps adaptively, producing sparse trajectories of $\sim 100$ points in length. We terminate adaptive time-stepping once the integrator either attempts to step to a point with $p<p_{\mathrm{stop}}$, where $p_{\mathrm{stop}}=p_{\mathrm{sep}}+\Delta p_{\mathrm{buf}}$ with $\Delta p_{\mathrm{buf}}=2\times 10^{-3}$, or exceeds the maximum duration $T$ set by the user [^8] Recall that we terminate our grid at $p_{\mathrm{sep}}+10^{-3}$, both to avoid numerical instabilities in the adiabatic equations of motion near the separatrix and because the dynamics in this regime are better modelled in a transition-to-plunge (rather than inspiral) framework. We then choose $\Delta p_{\mathrm{buf}}=2\times 10^{-3}$ to prevent trajectories from reaching this grid boundary.. In the former case, we must then place a point at $p=p_{\mathrm{stop}}$ to avoid truncation errors in the results of data analysis. As the adaptive stepping procedure first attempted to step past $p_{\mathrm{stop}}$, we have access to a continous solution in this interval; we therefore perform a numerical root-finding operation via Brent’s method to find the value of $t$ where
@@
is equal to zero. This typically converges to within $\sigma_{\mathrm{tol}}$ of the root in $\sim 5$ iterations and contributes little to the overall cost of a trajectory evaluation. Once the root $t_{\mathrm{root}}$ has been found, the final trajectory point is obtained by evaluating the continuous solution at $t_{\mathrm{root}}$. In the latter case of a trajectory duration exceeding $T$, this root-finding operation is not required and the continuous solution is simply evaluated at $t=T$ in order to obtain the last trajectory point.
@@ III.2.1 Teukolsky amplitude data grids
    The GW strain naturally decomposes onto a basis of spheroidal harmonics that, like the amplitudes, also evolve in time – note the dependence on $\hat{\omega}_{mn}$ in the spheroidal harmonic in Eq. 5.
To address the former issue, we simply precompute more waveform harmonics for the Kerr model, specifically all $(\ell,m,n)$ modes within the limits $102\leq\ell\leq\ell_{\mathrm{max}}=$ 10, $|m|\leq\ell$, and $|n|\leq n_{\mathrm{max}}=55$. For computational efficiency, we only compute mode amplitudes with $m\geq 0$ and infer amplitudes for $m<0$ with a conjugate mode symmetry Hughes:1999bq ; Fujita:2020zxe , as was performed in few v1 Katz:2021yft .
@@
where $\ell_{\mathrm{min}}=\mathrm{max}(2,|m|)$ and $b_{\ell mn}^{j}(t)$ are spherical-spheroidal mixing coefficients. Further discussion regarding this expansion [and the computation of $b_{\ell mn}^{j}(t)$] can be found in, e.g., Appendix A of Hughes:1999bq . The time dependence of the spheroidal harmonics is then absorbed by $b_{\ell mn}^{j}(t)$. Inserting Eq. 10 into Eq. 1, we then sum over the spheroidal harmonic mode amplitudes $A_{\ell mn}$ to obtain their corresponding spherical harmonic amplitudes, [^9] See also, e.g., Appendix A of Yunes:2010zj , Sec. IIIB of vandeMeent:2016pee and Sec. IIB of Nasipak:2023kuf for explicit derivations of this procedure.
@@
As Eq. 12 expresses the waveform strain for eccentric equatorial Kerr inspirals in the same form as that of Schwarzschild inspirals, we are free to apply few’s rapid mode summation framework without modification. This approach is readily extendable to eccentric and inclined Kerr inspirals once flux and mode amplitude data for these systems are available Fujita:2009us ; Fujita:2025pc ; Katz:2021yft ; Hughes:2021exa ; Chen:2023lsa ; Kerachian:2023oiw .
@@ III.2.2 Interpolation of mode amplitudes
Due to the large number of harmonic modes that contribute significantly to a typical EMRI waveform, obtaining their amplitudes efficiently via interpolation is a vital component of the few waveform generation framework. An important feature of this interpolation problem is that it is highly parallelisable with respect to both the mode indices $(\ell,m,n)$ and the parameters $(a,p,e)$. It is therefore sensible to design an interpolation framework that exploits this by computing mode amplitudes over many mode index combinations and parameter values simultaneously. For few v1, this was achieved by fitting a reduced-order model (in the form of a neural network) that rapidly evaluated a set of basis coefficients for multiple sets of input parameters. These coefficients were then converted into a complete set of mode amplitudes (i.e. over all sets of mode indices) via a matrix dot product; see Section 4 of Katz:2021yft for further discussion. This approach is computationally efficient and well-suited to acceleration via GPU vectorisation techniques. However, it suffers two significant drawbacks. First, the model struggles to accurately estimate the amplitudes of weaker (but still significant) modes. This was identified in previous work and was observed to limit the accuracy of the waveforms obtained to mismatches $\sim 10^{-5}$ Katz:2021yft , which is on the order of the accuracy required for unbiased analysis of the loudest expected astrophysical EMRI signals (which may have SNRs in the high hundreds Babak:2017tow ; see Section IV.3 for further discussion). Note that for quieter sources, larger mismatches than this will not necessarily lead to parameter biases Katz:2021yft ; such a requirement is only enforced by the (potential) existence of particularly loud (and therefore informative) signals.
@@
We instead opt to perform spline interpolation, which is particularly well-suited to this task and addresses the limitations outlined in the previous paragraph. Interpolating each mode amplitude separately ensures that the weaker modes in the spectrum are still estimated accurately. Mode amplitudes are readily interpolated in parallel, but as each interpolation operation is easily separable from the others, it is straightforward to evaluate only a requested sub-set of the full mode spectrum. The relationship between $\mathcal{A}_{\ell mn}$ and $a$ is sufficiently weak (at least for non-extremal systems burke2020constraining ; gralla2016inspiral ; Nasipak:2023kuf ) that linear interpolation with respect to this parameter performs adequately over the majority of the parameter space, further reducing computational costs. We therefore interpolate mode amplitude data with a “bicubic+linear” spline interpolation framework, striking a balance between accuracy and efficiency. Similarly to the case of the fluxes, we define data grid coordinates $(u,w,z)$ connected to $(p,e,a)$ by invertible transformations that are described in Appendix B. These transformations are of similar form to those used for the forcing function grid setup, but with adjusted coefficients to ensure the strong-field region is sufficiently densely sampled to capture the rapid amplitude variations that are characteristic of this region of the parameter space. The form of the eccentricity tapering is correspondingly altered to ensure it retains the same shape as the flux grids, such that the forcing function and amplitude interpolation domains (and therefore, the domain over which the waveform is defined) are consistent. The value of the mode amplitude $\mathcal{A}_{\ell mn}$ given grid coordinates $(u,w,z)$ is then written as
@@
where $\mathcal{A}_{\ell mn}^{\mathrm{-,+}}\equiv\mathcal{A}_{\ell mn}(\hat{Z}_{\mathrm{-,+}})$ are obtained via a bicubic spline interpolant over $(u,w)$, and $(\hat{Z}_{\mathrm{-}},\hat{Z}_{\mathrm{+}})$ are the values of the grid points immediately (below, above) $z$.
@@
[…]

### v2

@@ III.1.2 Rapid computation of ODE derivatives
In order to integrate Eqs.˜2 and 3, we must first specify the form of $\hat{f}^{(0)}_{p,e}$ and $\hat{\Omega}_{\phi,r}$. The latter terms are simple to incorporate; analytic expressions for the dimensionless orbital frequencies of geodesics in generic Kerr, which are equal to $\hat{\Omega}_{\phi,r}$ at adiabatic order, have existed in the literature for some time (see e.g. Fujita:2009bp ; vandeMeent:2019cam and references therein) and are inexpensive to evaluate. We employ these expressions in our framework, simplifying terms involving $x_{I}$ (given $x_{I}=1$) where possible for efficiency and numerical stability.
Obtaining the forcing functions is less straightforward, as the data generation procedure in Section˜III.1.1 is orders of magnitude too expensive to be applied on-the-fly during trajectory integration. However, these fluxes vary smoothly over the parameter space and are ideal targets for interpolation schemes. We therefore opt to generate forcing function values en-masse, producing data grids in parameter space over which to perform this interpolation. Producing these data grids is computationally expensive, but must only be performed once; as the methods described in Section˜III.1.1 are straightforward to parallelise, this can be achieved relatively quickly on distributed computing resources. Of particular note is the accuracy to which the fluxes must be estimated, which is $\mathcal{O}(\nu)$ for sub-radian accumulated phase error over the course of a typical trajectory evaluation.
@@
Next, we must specify our interpolation variables. This is essential for our framework because a regular grid cannot be constructed in $(a,p,e)$ space (recalling that $x_{I}=1$ for equatorial inspirals) due to the boundary imposed by $p_{\mathrm{sep}}$ that varies across the parameter space. Choosing sensible interpolation variables can also spread out rapid variation in the fluxes across the parameter space (especially near the separatrix), improving the accuracy of interpolation in these regions. To this end, we format our data grids in terms of three interpolation variables $(u,w,z)$ that are related to $(a,p,e)$ by bijective and analytically-invertible transformations. For brevity, we will not state these transformations in the main text here (they are given in full in Appendix˜B) but instead summarize the important features of the resulting grid structure below:
@@
The shape and grid-point distribution of these grids is also shown in Fig.˜20 and Fig.˜21 in Appendix˜B, which clearly displays the structure of the eccentricity taper. Functions for performing these mappings are implemented in few and are readily applicable in the construction of user-specified grids.
Finally, we must choose the dimensions of the grid. With all other aspects of the problem fixed, this is essentially a trade-off between the accuracy of the interpolant and the computational cost of producing the grids. However, assessing how interpolation error in trajectory modelling propagates to the accuracy of the resulting waveform is not straightforward. We therefore opt to identify empirically what flux grid dimensions are required to satisfy waveform accuracy requirements. Based on this process, we settle on the dimensions $(N_{u},N_{w},N_{z})$ = $(129,65,65),(65,33,33)$ for the inner and outer grid regions respectively. As we will later demonstrate in Section˜IV, this ensures that grid point density (and therefore, the accuracy of the interpolated forcing functions) is sufficient for our model to be highly robust over the majority of the parameter space.
@@ III.1.3 ODE integration with continuous solution
A continuous approximation for the phase also yields similar piecewise polynomials for the orbital frequencies and their derivatives with respect to time; as shown in Fig.˜2, this approximation is very accurate (exceeding 9 decimal places in the frequencies for the majority of the inspiral). This is particularly useful in the construction of frequency-domain waveforms (which require this information as an input) and will enable resonance effects (which require careful monitoring of these frequencies, as well as accurate frequency derivatives) to be seamlessly folded into the few framework in the future.
@@ III.1.4 Numerical integration of ODE system
With our framework for obtaining the derivatives in Eqs.˜2 and 3 and integrating the resulting ODE system in place, we are ready to evolve a trajectory from a set of initial conditions $\{p_{0},e_{0},\Phi_{\phi 0},\Phi_{r0}\}$ given specific values for $M$, $\nu$ and $a$. To improve the numerical stability of the integrator, we first rescale the problem such that the integration variables vary by $\mathcal{O}(1)$ across an inspiral. We achieve this by integrating the system with respect to the radiation-reaction timescale $t_{\mathrm{rr}}=\nu t$ and eliminating the pre-factor $\nu$ from Eq.˜2 accordingly. Rather than scaling Eq.˜3 similarly, we leave it unchanged (in order to achieve the desired property of $\mathcal{O}(1)$ scaling for all parameters) and instead rescale the orbital phase evolution post-integration by $\nu^{-1}$. This approach is valid in the adiabatic approximation because (by definition) the $\hat{f}^{(0)}_{p,e}$ are independent of the orbital phases. A consequence of this choice is that, for an integrator absolute error tolerance $\sigma_{\mathrm{tol}}$, the effective error tolerance on the orbital phases is $\sigma_{\mathrm{tol}}/\nu$. This is in fact a desirable outcome: we are mainly interested in tuning the accuracy with which $p(t)$ and $e(t)$ are evolved (as small errors in these parameters will gradually accumulate over an inspiral, coupling in turn to the orbital phase evolution), whereas the error in the reconstructed phase will remain of order $\sigma_{\mathrm{tol}}/\nu$, which is small and has negligible impact on waveform phase accuracy. This scaling is demonstrated in Fig.˜3, where phase error grows in proportion to (but is several orders of magnitude larger than) the error tolerance of the ODE integrator. Trajectory wall-time varies significantly with respect to ODE error; to ensure trajectories are accurate to within $10^{-3}$ radians, we set $\sigma_{\mathrm{tol}}=10^{-11}$ by default, but this can be relaxed by the end user according to their accuracy requirements. We also note that trajectory wall-time will be hardware-dependent; the timings presented in Fig.˜3 were obtained with an Apple M3 processor.
@@
Trajectory integration can be performed either forwards or backwards with respect to time (the latter achieved by negating Eqs.˜2 and 3). The size of the first integration step from the starting point is chosen automatically with standard techniques GLADWELL1987175 . The integrator then steps adaptively, producing sparse trajectories of $\sim 100$ points in length. During forwards integration, we terminate adaptive time-stepping once the integrator either attempts to step to a point with $p<p_{\mathrm{stop}}$, where $p_{\mathrm{stop}}=p_{\mathrm{sep}}+\Delta p_{\mathrm{buf}}$ with $\Delta p_{\mathrm{buf}}=2\times 10^{-3}$, or exceeds the maximum duration $T$ set by the user [^8] Recall that we terminate our grid at $p_{\mathrm{sep}}+10^{-3}$, both to avoid numerical instabilities in the adiabatic equations of motion near the separatrix and because the dynamics in this regime are better modelled in a transition-to-plunge (rather than inspiral) framework. We then choose $\Delta p_{\mathrm{buf}}=2\times 10^{-3}$ to prevent trajectories from reaching this grid boundary.. In the former case, we must then place a point at $p=p_{\mathrm{stop}}$ to avoid truncation errors in the results of data analysis. As the adaptive stepping procedure first attempted to step past $p_{\mathrm{stop}}$, we have access to a continous solution in this interval; we therefore perform a numerical root-finding operation via Brent’s method to find the value of $t$ where
@@
is equal to zero. This typically converges to within $\sigma_{\mathrm{tol}}$ of the root in $\sim 5$ iterations and contributes little to the overall cost of a trajectory evaluation. Once the root $t_{\mathrm{root}}$ has been found, the final trajectory point is obtained by evaluating the continuous solution at $t_{\mathrm{root}}$. In the latter case of a trajectory duration exceeding $T$, this root-finding operation is not required and the continuous solution is simply evaluated at $t=T$ in order to obtain the last trajectory point. Backwards integration proceeds similarly, except that $p_{\mathrm{stop}}$ is now determined by the boundaries of the forcing function data grids (see Appendix˜B for these definitions). If backwards integration reaches this grid boundary before the duration exceeds $T$, a final point is placed at the boundary and integration is terminated.
@@ III.2.1 Teukolsky amplitude data grids
    The GW strain naturally decomposes onto a basis of spheroidal harmonics that, like the amplitudes, also evolve in time – note the dependence on $\hat{\omega}_{mn}$ in the spheroidal harmonic in Eq.˜5.
To address the former issue, we simply precompute more waveform harmonics for the Kerr model, specifically all $(\ell,m,n)$ modes within the limits $2\leq\ell\leq\ell_{\mathrm{max}}=10$, $|m|\leq\ell$, and $|n|\leq n_{\mathrm{max}}=55$. For computational efficiency, we only compute mode amplitudes with $m\geq 0$ and infer amplitudes for $m<0$ with a conjugate mode symmetry Hughes:1999bq ; Fujita:2020zxe , as was performed in few v1 Katz:2021yft .
@@
where $\ell_{\mathrm{min}}=\mathrm{max}(2,|m|)$ and $b_{\ell mn}^{j}(t)$ are spherical-spheroidal mixing coefficients. Further discussion regarding this expansion [and the computation of $b_{\ell mn}^{j}(t)$] can be found in, e.g., Appendix A of Hughes:1999bq . The time dependence of the spheroidal harmonics is then absorbed by $b_{\ell mn}^{j}(t)$. Inserting Eq.˜10 into Eq.˜1, we then sum over the spheroidal harmonic mode amplitudes $A_{\ell mn}$ to obtain their corresponding spherical harmonic amplitudes, [^9] See also, e.g., Appendix A of Yunes:2010zj , Sec. IIIB of vandeMeent:2016pee and Sec. IIB of Nasipak:2023kuf for explicit derivations of this procedure.
@@
As Eq.˜12 expresses the waveform strain for eccentric equatorial Kerr inspirals in the same form as that of Schwarzschild inspirals, we are free to apply few’s rapid mode summation framework without modification. This approach is readily extendable to eccentric and inclined Kerr inspirals once flux and mode amplitude data for these systems are available Fujita:2009us ; Fujita:2025pc ; Katz:2021yft ; Hughes:2021exa ; Chen:2023lsa ; Kerachian:2023oiw .
@@ III.2.2 Interpolation of mode amplitudes
Due to the large number of harmonic modes that contribute significantly to a typical EMRI waveform, obtaining their amplitudes efficiently via interpolation is a vital component of the few waveform generation framework. As was performed in few v1, we opt to interpolate the complex amplitudes in terms of their real and imaginary components. These quantities vary smoothly over the parameter space, and this allows for both the leading-order term $|\mathcal{A}_{\ell mn}|$ and the post-adiabatic contribution $\psi_{\ell mn}=\arg(\mathcal{A}_{\ell mn})$ to be obtained directly, at the cost of the additional memory needed to store data grids for both components at full resolution. While the latter term is not strictly-speaking necessary for an adiabatic model, we seek to demonstrate that all effects at first post-adiabatic order can readily be included in the FEW framework, and so include this for completeness. While it is tempting to interpolate $|\mathcal{A}_{\ell mn}|$ and $\psi_{\ell mn}$ directly (with the accuracy requirements on the latter being relaxed due to its post-adiabatic nature), the former is non-smooth (due to zero-crossings in the amplitude surfaces) and the latter must be unwrapped to avoid discontinuities (which is challenging in multiple dimensions). We plan to investigate whether a form of amplitudes exists that retains the benefits of both conventions in more detail in future work.
An important feature of this interpolation problem is that it is highly parallelisable with respect to both the mode indices $(\ell,m,n)$ and the parameters $(a,p,e)$. It is therefore sensible to design an interpolation framework that exploits this by computing mode amplitudes over many mode index combinations and parameter values simultaneously. For few v1, this was achieved by fitting a reduced-order model (in the form of a neural network) that rapidly evaluated a set of basis coefficients for multiple sets of input parameters. These coefficients were then converted into a complete set of mode amplitudes (i.e. over all sets of mode indices) via a matrix dot product; see Section 4 of Katz:2021yft for further discussion. This approach is computationally efficient and well-suited to acceleration via GPU vectorisation techniques. However, it suffers two significant drawbacks. First, the model struggles to accurately estimate the amplitudes of weaker (but still significant) modes. This was identified in previous work and was observed to limit the accuracy of the waveforms obtained to mismatches $\sim 10^{-5}$ Katz:2021yft , which is on the order of the accuracy required for unbiased analysis of the loudest expected astrophysical EMRI signals (which may have SNRs in the high hundreds Babak:2017tow ; see Section˜IV.3 for further discussion). Note that for quieter sources, larger mismatches than this will not necessarily lead to parameter biases Katz:2021yft ; such a requirement is only enforced by the (potential) existence of particularly loud (and therefore informative) signals.
@@
We instead opt to perform spline interpolation, which is particularly well-suited to this task and addresses the limitations outlined in the previous paragraph. Interpolating each mode amplitude separately ensures that the weaker modes in the spectrum are still estimated accurately. Mode amplitudes are readily interpolated in parallel, but as each interpolation operation is easily separable from the others, it is straightforward to evaluate only a requested sub-set of the full mode spectrum. The relationship between $\mathcal{A}_{\ell mn}$ and $a$ is sufficiently weak (at least for non-extremal systems burke2020constraining ; gralla2016inspiral ; Nasipak:2023kuf ) that linear interpolation with respect to this parameter performs adequately over the majority of the parameter space, further reducing computational costs. We therefore interpolate mode amplitude data with a “bicubic+linear” spline interpolation framework, striking a balance between accuracy and efficiency. Similarly to the case of the fluxes, we define data grid coordinates $(u,w,z)$ connected to $(p,e,a)$ by invertible transformations that are described in Appendix˜B. These transformations are of similar form to those used for the forcing function grid setup, but with adjusted coefficients to ensure the strong-field region is sufficiently densely sampled to capture the rapid amplitude variations that are characteristic of this region of the parameter space. The form of the eccentricity tapering is correspondingly altered to ensure it retains the same shape as the flux grids, such that the forcing function and amplitude interpolation domains (and therefore, the domain over which the waveform is defined) are consistent. The value of the mode amplitude $\mathcal{A}_{\ell mn}$ given grid coordinates $(u,w,z)$ is then written as
@@
To justify the wide range of modes we include in our waveform model, in Fig.˜4 we apply our amplitude interpolant to demonstrate how the morphology of the waveform mode spectrum changes over an eccentric $(e_{0}=0.8)$ inspiral into a rapidly-spinning $(a=0.998$) MBH, plotting the mode power $P_{\ell mn}=|A_{\ell mn}|^{2}$ as a function of $(\ell,m,n)$. At early times, the waveform contains rich $n$-mode content due to the high initial eccentricity, but the mode spectrum rapidly tails off with respect to $\ell$ as the CO is orbiting in the relatively weak field. As the inspiral circularizes due to GW emission, the higher-order $n$-mode contribution gradually weakens; meanwhile, the high spin of the MBH allows the CO trajectory to extend deep into the strong field of the MBH, significantly accentuating the higher $\ell$-mode content of the waveform. Typically, for systems with higher eccentricities at lower semi-latus rectum values than what is shown in Fig.˜4, more power is contained in the higher $n$-mode content of the spectrum (particularly for higher values of $\ell$) Drasco:2005kz ; Fujita:2009us .
One significant drawback of interpolating each mode amplitude with cubic splines (as opposed to a reduced-basis representation) is that the size of the spline coefficient arrays that must be stored in memory grows linearly with the number of modes supported in the model. For the 6993 modes that we interpolate in our waveform model, these coefficients arrays total $\sim$5\text{\,}\mathrm{GB}$\$ in size, which is sufficiently compact to fit in the memory of most GPUs. However, some care will be required once few is extended to inclined and eccentric inspirals in future work, as the number of waveform modes (and therefore the memory requirements) will increase significantly (with typically as many as $\sim 10^{5}-10^{6}$ waveform modes Katz:2021yft ; Hughes:2021exa ; Isoyama:2021jjd ).